The particular integral value of (D² + 4)y = sin²x is


A) 1/8-x/8 sin2x


B) 1/8+x/8 Sin2x


C) 1/8-x/8 Cos2x D)1/8+x/8 Cos2x


1
Expert's answer
2021-08-20T01:36:57-0400

"(D^2+4)y=sin^2x"

complementary function (C.F.) is obtained by putting

"D^2+4=0\u27f9D=\u00b12i"

"\u2234C.F.=C_1cos2x+C_2sin2x"

Now, a particular integral (P.I.) is found out as follows

"P.I.=\frac{sin^2x}{D^2+4}\\
=\frac{\frac{1−cos2x}{2}}{D2+4}\\
=\frac{1}{2}(\frac{1−cos2x}{D^2+4})\\
=\frac{1}{2}(\frac{1}{D^2+4}−\frac{cos2x}{D^2+4})\\
=\frac{1}{2}(\frac{1}{0^2+4}−\frac{x}{2} \frac{1}{D}(cos2x))\\
=\frac{1}{2}(\frac{1}{4}−\frac{x}{2}\frac{sin2x}{2})\\
=\frac18−\frac{xsin2x}{8}"